# Alljoined - A dataset for EEG-to-Image decoding

### v1

@@ Abstract
We present Alljoined, a dataset built specifically for EEG-to-Image decoding. Recognizing that an extensive and unbiased sampling of neural responses to visual stimuli is crucial for image reconstruction efforts, we collected data from 8 participants looking at 10,000 natural images each. We have currently gathered 46,080 epochs of brain responses recorded with a 64-channel EEG headset. The dataset combines response-based stimulus timing, repetition between blocks and sessions, and diverse image classes with the goal of improving signal quality. For transparency, we also provide data quality scores. We publicly release the dataset and all code at https://linktr.ee/alljoined1.
* equal contribution. core contribution. ${}^{1}$llJoined ${}^{2}$he University of Waterloo ${}^{3}$niversité Paris-Saclay, Inria TAU, CNRS, LISN ${}^{4}$ederal University of ABC ${}^{5}$rije Universiteit Amsterdam ${}^{6}$niversity of Toronto
@@ 1 Introduction
In the fields of cognitive neuroscience and medical imaging, advancements in deep learning have demonstrated unparalleled precision in decoding brain activity [36, 20, 38, 37, 6]. Researchers have translated the intricate patterns of brain activity during various cognitive processes by utilizing neuroimaging modalities, such as functional Magnetic Resonance Imaging (fMRI) and electroencephalography (EEG).
@@
While fMRI has traditionally been the primary tool for image reconstruction due to its excellent spatial resolution, the slow temporal components severely delimit actual clinical usage. On the other hand, EEG is a medical modality available in everyday clinical contexts with an excellent time resolution [33, 40, 39]. As neurons fire at millisecond scales, the high temporal resolution provided by EEG is crucial for real-time monitoring of neural dynamics [45, 12, 17]. Additionally, EEG is portable, more accessible to set up, and much more cost-effective than fMRI, making it suitable for real-world applications, including brain-computer interfaces and clinical diagnostics.
The development of very large fMRI-to-image datasets has proven foundational for recent breakthroughs in deep-learning image reconstruction projects. Inspired by the need for such datasets in the EEG domain, we present Alljoined, a novel, large-scale dataset covering a wide range of naturalistic stimuli that allows for robust, generalizable image reconstruction efforts. Our contributions are as follows:
@@ 2.1 EEG-to-Image Datasets
A popular EEG-image dataset is Brain2Image [22], which consists of evoked responses to a visual stimulus from distinct image classes. Each block consists of stimuli corresponding only to a single image class. There are 40 classes, with 50 unique images in each class. This dataset has been criticized for having no train-test separation during recording, block-specific stimuli patterns, and lack of consistency across different frequency bands. These factors can incorrectly boost model performance by giving extraneous proxy information about the block rather than the actual image-specific brain responses [28, 3]. Moreover, the limited number of classes may encourage models to perform 40-way classification rather than image reconstruction.
@@
Recent studies achieving impressive reconstruction results have relied on this dataset [26, 5, 23], which may affect the validity of their results. As recommended by [28, 3], the stimuli within each block in our dataset were chosen randomly across a variety of natural images, effectively minimizing the risk for block-class correlations.
@@
The diversity of decoding stimuli further limits current EEG-based image reconstruction datasets. While Brain2Image consists of 40 classes, several studies utilize a dataset of visual imagery of characters and objects belonging to only 10 different classes, ThoughtViz [46]. Both Brain2Image and ThoughtViz fail to represent the continuous, diverse quality of naturalistic stimuli. The same limitation applies to studies utilizing a severely limited quantity of naturalistic stimuli. Approaches to EEG-based image reconstruction derived from the ThoughtViz [32, 39], Brain2Image [26, 5, 23], or other equally selective datasets [2, 48], may thus suffer from generalizing well to diverse, real-world stimuli.
@@
To account for the diverse and continuous nature of naturalistic images, Alljoined consists of 1) 10,000 images per participant and 2) that belong to at least one of 80 MS-COCO [30] object categories. Importantly, each MS-COCO category is broader than a single object class (e.g. the things category includes car, skateboard, hat, etc.), and each image can belong to up to 5 classes [29].
@@
There are also existing datasets that include naturalistic stimuli, but compromise in other domains. The MindBigData initiative [47] captures a wide sample of images from the ImageNet dataset, but is derived only from a single individual, limiting the potential of training image reconstruction models that generalize to other individuals. The THINGS-EEG1 [16] and THINGS-EEG2 [13] datasets were acquired using short image presentation times of 50 and 100 ms, and a stimulus onset asynchrony of 100 and 200 ms.
@@
Although the rapid serial visual presentation (RSVP) [15] paradigm proposes disentangling the temporal dynamics of visual processing and categorical abstraction of non-target stimuli, it is not ideal for capturing cortical image processing beyond early visual activity with low noise. We see that [41] obtained the highest accuracy with their EEG-image classifier when focusing on 320-480 ms after stimulus onset, and [34] is able to extract relevant decoding features even around 550 ms after stimulus onset. This suggests that while it takes 50-120 ms for object recognition of a stimulus to register in the visual cortex, a longer stimulus period is beneficial for accuracy on downstream tasks. Alljoined consists of extensive data from n=8 participants, measured with an inter-stimulus interval of 300 ms, which captures important hallmarks in visual processing while maintaining a high presentation frequency [44, 16].
@@ 3.1 Participants
We collected data from eight participants (six male, two female), with an average age of 22 $\pm$ 0.64 years, all with normal or corrected-to-normal vision, right-handed. All participants were healthy, with no neurocognitive impairments, except 2 participants who reported a history of mental health disorders (e.g. GAD, ADHD). Each participant provided informed consent. The Research Ethics Board approved the procedures as suppressed for double-blind review.
@@ 3.3 Procedure
Images were displayed to participants over the course of multiple one-hour-long sessions. Each session consisted of 16 blocks, wherein images in the first 8 unique blocks were in the second 8 blocks.
@@ 3.5 Pre-processing
Initially, we applied the band-pass filtering with a low frequency of $0.5$ Hz and a high frequency of $125$ Hz with overlap-add finite impulse response filtering, with range based on [43]. We then apply a notch filter at 60Hz to eliminate power line noise.
@@
We segment our data into intervals starting at $-50$ ms onset stimulus and ending at the end of each trial at $600$ ms, as the stimulus section describes in Section 3.2. We discard the jitter time between trials.
@@
Finally, we re-reference our channels using an average reference scheme, before applying a baseline correction window from $-50$ ms to 00 ms relative to stimulus onset, following [42] recommendation for ERP baseline. The epoch data subtracts the average activation during the baseline interval to remove noise from the signal.
@@ 4 Analysis
Figure 4: EEG topographic maps and corresponding signals at all 64 electrodes averaged over a) 3823 events for the fifth participant (left) and b) across all sessions for all participants (43070 events) in the Alljoined1 dataset (right), highlighting individual and common brain activity patterns associated with image presentation.
@@ 4.1 ERP Analysis
Topographies of activation additionally reveal a strong concentration of positive activation at occipital, parietal, and partly temporal electrodes and a consistently negative activation at central and frontal areas. This topographical distribution was stable across the duration of the ERP and corresponded well between the single participant and the cohort. Given the strong peak in activity at the occipital and parietal areas, we further investigated the distribution of ERPs across individual participants and sessions at the occipital and parietal electrodes, as displayed in figure 5. While the magnitude of activation differs between participants, we conclude a by-and-large consistent activation pattern across participants and sessions.
@@ 4.2 SNR Analysis
The Signal-to-Noise Ratio (SNR) serves as a pivotal metric in evaluating the efficacy of our dataset. To ascertain the SNR, we employ the Standard Measurement Error (SME) as a gauge for noise assessment [31]. The SME is determined by calculating the standard deviation of the aggregated waveform average for each event type across all trials and then dividing this by the square root of the event type’s occurrence count. The SNR is subsequently derived by dividing the mean signal values by their corresponding SME.
Figure 6 compares the average SNR across all events in a single session for participant 5 with the average SNR across all events for both sessions concatenated. We see that the SNR is noticeably lower in the multi session graph. This is due to the increased number of repetitions for a given event at different timepoints. This attributes to a disproportionately higher standard deviation value and consequently a higher SME and lower SNR. However, that is not to say that the quality of the data is worse. It actually reflects more accurate SNR values as there are more data points, distributed across different sessions. It is also observed that the single sessions graph is more volatile across time, demonstrating a greater variance in SNR values which are captured by have a less accurate metric for noise with less trials to average between. This fluctuation underscores the limited accuracy of noise metrics derived from fewer trials, thus highlighting the critical importance of incorporating repeated measures across sessions or blocks for robust SNR evaluation.
@@ 4.3 Discussion
In conclusion, the design parameters of our experiment not only accommodate the full breadth of neural responses captured in the critical window of 300ms but also provide a comparative advantage over the shorter intervals used in THINGS EEG2 and THINGS EEG1. The ideal timing of our experiment ensures the acquisition of a comprehensive ERP waveform, contributing to a more nuanced understanding of cognitive processes and neural dynamics.
@@ 5 Conclusion
We introduce Alljoined1, an EEG-image dataset that uses well-timed stimuli, repetitions between blocks and sessions, and a wide distribution of natural images to create an improved dataset for image decoding tasks. We believe that its size, diversity, and quality will help promote work to better understand the mechanisms of visual processing, and in decoding visual responses in clinical and consumer BCI contexts.
Future Directions: We are eager to explore high-density EEG recording of exclusively the occipital and parietal regions to better target regions of the brain most responsive to visual stimuli. We are also interested in conducting ablation studies on the generalizability of responses to imagined mental imagery. We also believe there is great potential in exploring continuous data collection in natural environments with a wireless headset.
Data availability: The preprocessed EEG dataset is available on OSF. Labels to the corresponding NSD image IDs are included in the object files.
Code availability: The stimulus and preprocessing code to reproduce all the results is available on anonymized GitHub here and here.
@@ 6 Acknowledgements
This work was sponsored by ZFellows. BA work was supported by DATAIA Convergence Institute as part of the “Programme d’Investissement d’Avenir”, (ANR-17-CONV-0003) operated by LISN-CNRS. We would like to thank Dr Sylvain Chevallier for his valuable feedback on this manuscript.

### v3

# Alljoined1 - A dataset for EEG-to-Image decoding

@@ Abstract
We present Alljoined1, a dataset built specifically for EEG-to-Image decoding. Recognizing that an extensive and unbiased sampling of neural responses to visual stimuli is crucial for image reconstruction efforts, we collected data from 8 participants looking at 10,000 natural images each. We have currently gathered 46,080 epochs of brain responses recorded with a 64-channel EEG headset. The dataset combines response-based stimulus timing, repetition between blocks and sessions, and diverse image classes with the goal of improving signal quality. For transparency, we also provide data quality scores. We publicly release the dataset and all code at https://linktr.ee/alljoined1.
* equal contribution. core contribution. 1Alljoined 2The University of Waterloo 3Université Paris-Saclay, Inria TAU, CNRS, LISN 4Federal University of ABC 5Vrije Universiteit Amsterdam 6University of Toronto
@@ 1 Introduction
In the fields of cognitive neuroscience and medical imaging, advancements in deep learning have led to unparalleled precision in decoding brain activity [38, 21, 40, 39, 7]. Researchers have translated the intricate patterns of brain activity during various cognitive processes by utilizing neuroimaging modalities, such as functional Magnetic Resonance Imaging (fMRI) and electroencephalography (EEG).
@@
While fMRI has traditionally been the primary tool for image reconstruction due to its excellent spatial resolution, its low temporal resolution severely delimits actual clinical usage. On the other hand, EEG is a medical modality available in everyday clinical contexts with an excellent time resolution [35, 42, 41]. As neurons fire at millisecond scales, the high temporal resolution provided by EEG is crucial for real-time monitoring of neural dynamics [47, 13, 18]. Additionally, EEG is portable, more accessible to set up, and much more cost-effective than fMRI, making it suitable for real-world applications, including brain-computer interfaces and clinical diagnostics.
The development of very large fMRI-to-image datasets has proven foundational for recent breakthroughs in deep-learning image reconstruction projects. Inspired by the need for such datasets in the EEG domain, we present Alljoined1, a novel, large-scale dataset covering a wide range of naturalistic stimuli that allows for robust, generalizable image reconstruction efforts. Our contributions are as follows:
@@ 2.1 EEG-to-Image Datasets
A popular EEG-image dataset is Brain2Image [23], which consists of evoked responses to a visual stimulus from distinct image classes. Each block consists of stimuli corresponding only to a single image class. There are 40 classes, with 50 unique images in each class. This dataset has been criticized for having no train-test separation during recording, block-specific stimuli patterns, and lack of consistency across different frequency bands. These factors can incorrectly boost model performance by giving extraneous proxy information about the block rather than the actual image-specific brain responses [29, 4]. An extensive study highlights how many recent EEG-based image reconstruction attempts depend crucially on their block design, demonstrating how similar analytical approaches are not capable of meaningfully decoding EEG signals in a rapid serial visual paradigm (RSVP) [29], even when collecting large amounts of data for only a single subject [3].
@@
Recent studies achieving impressive reconstruction results have relied on data collected with flawed block designs [27, 6, 24], calling the validity of their results into question. As recommended by [29, 4], the stimuli within each block in our dataset were chosen randomly across a variety of natural images, effectively minimizing the risk of block-class correlations.
@@
The diversity of decoding stimuli further limits current EEG-based image reconstruction datasets. While studies like Brain2Image or [3] consist of images belonging to 40 classes, several studies utilize a dataset of visual imagery of characters and objects belonging to only 10 different classes, ThoughtViz [48]. Such a discretized representation of real world objects fails to account for the continuous, diverse quality of naturalistic stimuli. The same limitation applies to studies utilizing a severely limited quantity of naturalistic stimuli. Approaches to EEG-based image reconstruction derived from the ThoughtViz [34, 41], Brain2Image [27, 6, 24], or other equally selective datasets [2, 50], may thus suffer from generalizing well to diverse, real-world stimuli. Moreover, a limited number of image classes may encourage image reconstruction models to generate images class-conditionally, rather than reconstructing images based on (continuous) brain-encoded semantic or perceptual attributes of an image.
@@
To account for the diverse and continuous nature of naturalistic images, Alljoined1 consists of 1) 10,000 images per participant 2) that belong to at least one of 80 MS-COCO [31] object categories. Importantly, each MS-COCO category is broader than a single object class (e.g. the things category includes car, skateboard, hat, etc.), and each image can belong to up to 5 classes [30].
@@
There are also existing datasets that include naturalistic stimuli, but compromise in other domains. The MindBigData initiative [49], or [3] capture a wide sample of images, but are derived only from a single individual, limiting the potential of training image reconstruction models that generalize to other individuals. The THINGS-EEG1 [17] and THINGS-EEG2 [14] datasets were acquired using short image presentation times of 50 and 100 ms, and a stimulus onset asynchrony of 100 and 200 ms.
@@
Although the rapid serial visual presentation [16] paradigm proposes disentangling the temporal dynamics of visual processing and categorical abstraction of non-target stimuli, it is not ideal for capturing cortical image processing beyond early visual activity with low noise. We see that [43] obtained the highest accuracy with their EEG-image classifier when focusing on 320-480 ms after stimulus onset, and [36] is able to extract relevant decoding features even around 550 ms after stimulus onset. This suggests that while it takes 50-120 ms for object recognition of a stimulus to register in the visual cortex, a longer stimulus period is beneficial for accuracy on downstream tasks. Alljoined1 consists of extensive data from eight participants, measured with an inter-stimulus interval of 300 ms, which captures important hallmarks in visual processing while maintaining a high presentation frequency [46, 17]. This setup might furthermore allow us to overcome the limitations in decoding image content from EEG activity in RSVP paradigms, as previously reported in [3].
@@ 3.1 Participants
We collected data from eight participants (six male, two female), with an average age of 22 $\pm$ 0.64 years, all with normal or corrected-to-normal vision, right-handed. All participants were healthy, with no neurocognitive impairments, except 2 participants who reported a history of mental health disorders (e.g. GAD, ADHD). Each participant provided informed consent. The Research Ethics Board approved the procedures as suppressed for double-blind review. We note that there are potential limitations of the study due to the imbalance between the genders of the participants and the low age disparity, which could influence bias and learning with AI models.
@@ 3.3 Procedure
Images were displayed to participants over the course of multiple one-hour-long sessions. Each session consisted of 16 blocks, wherein images in the first 8 unique blocks were repeated in the second 8 blocks.
@@ 3.5 Pre-processing
Initially, we applied band-pass filtering with a low frequency of $0.5$ Hz and a high frequency of $125$ Hz with overlap-add finite impulse response filtering, with range based on [45]. We then apply a notch filter at 60Hz to eliminate power line noise.
@@
We segment our data into intervals starting at $-50$ ms onset stimulus and ending at the end of each trial at $600$ ms, as the stimulus section describes in Section 3.2. Each segment of data is termed an epoch. We discard the jitter time between trials.
@@
Finally, we re-reference our channels using an average reference scheme, before applying a baseline correction window from $-50$ ms to 00 ms relative to stimulus onset, following recommendations from [44] for ERP baseline. The epoch data subtracts the average activation during the baseline interval to remove noise from the signal.
@@ 4 Analysis
Figure 4: EEG topographic maps and corresponding signals at all 64 electrodes averaged over a) 3823 events for the fifth participant (left) and b) across all sessions for all participants (43070 events) in the Alljoined1 dataset (right), highlighting individual and common brain activity patterns associated with image presentation. An event is defined as a specific time point in the experiment.
@@ 4.1 ERP Analysis
Topographies of activation additionally reveal a strong concentration of positive activation at occipital, parietal, and partially temporal electrode locations and a consistently negative activation at central and frontal areas. This topographical distribution was stable across the duration of the ERP and corresponded well between the single participant and the cohort. Given the strong peak in activity at the occipital and parietal areas, we further investigated the distribution of ERPs across individual participants and sessions at the occipital and parietal electrodes, as displayed in figure 5. While the magnitude of activation differs between participants, we conclude a by-and-large consistent activation pattern across participants and sessions.
@@ 4.2 SNR Analysis
The Signal-to-Noise Ratio (SNR) serves as a pivotal metric in evaluating the efficacy of our dataset. To ascertain the SNR, we employ the Standardized Measurement Error (SME) as a gauge for noise assessment [33]. We choose SME as our metric of choice as it is able to robustly quantify the data quality for each participant at each electrode site [32]. The SME is determined by calculating the standard deviation of the aggregated waveform average for each event type across all trials and then dividing this by the square root of the event type’s occurrence count. The SNR is subsequently derived by dividing the mean signal values by their corresponding SME.
Figure 6 compares the average SNR across all events in a single session for participant 5 with the average SNR across all events for both sessions concatenated. We see that the SNR is noticeably lower in the multi session graph. This is due to the increased number of repetitions for a given event at different timepoints. This leads to a disproportionately higher standard deviation value and consequently a higher SME and lower SNR. However, that is not to say that the quality of the data is worse. It actually reflects more accurate SNR values as there are more data points, distributed across different sessions. It is also observed that the single sessions graph is more volatile across time, demonstrating a greater variance in SNR values which are captured by having a less accurate metric for noise with less trials to average between. This fluctuation underscores the limited accuracy of noise metrics derived from fewer trials, thus highlighting the critical importance of incorporating repeated measures across sessions or blocks for robust SNR evaluation.
@@ 4.3 Discussion
In conclusion, we choose a 300ms latency as it provides a good trade-off between capturing long-term neural activity whilst maintaining a high presentation frequency. The ideal timing of our experiment ensures the acquisition of a comprehensive ERP waveform, contributing to a more nuanced understanding of cognitive processes and neural dynamics as compared to the shorter intervals used in THINGS EEG2 and THINGS EEG1.
@@ 5 Conclusion
We introduce Alljoined1, an EEG-image dataset that uses well-timed stimuli, repetitions between blocks and sessions, and a wide distribution of natural images to create an improved dataset for image decoding tasks. We believe that its size, diversity, and quality will help promote work to better understand the mechanisms of visual processing, and in decoding visual responses in clinical and consumer brain-computer interface (BCI) contexts.
Future directions: We are eager to explore high-density EEG recordings of exclusively the occipital and parietal regions to better target regions of the brain most responsive to visual stimuli. We are also interested in conducting ablation studies on the generalizability of responses to imagined mental imagery. We further believe there is great potential in exploring continuous data collection in natural environments with a wireless headset.
Data availability: Both the raw and preprocessed EEG dataset is available on OSF. Labels to the corresponding NSD image IDs are included in the object files.
Code availability: The stimulus and preprocessing code to reproduce all the results is available on Anonymous GitHub here and here.
@@ 6 Acknowledgements
This work was sponsored by Z Fellows, Hack Grants, Moth Fund, Fiona Leng and . BA work was supported by DATAIA Convergence Institute as part of the “Programme d’Investissement d’Avenir”, (ANR-17-CONV-0003) operated by LISN-CNRS. We would like to thank Dr Sylvain Chevallier for his valuable feedback on this manuscript.